OpenStudy (anonymous):

Suppose a culture of bacteria starts with 9000 bacteria. After one hour, the count is 10,000. Find an exponential equation that models the number of bacteria in hours. Keep at least 4 decimal places in your formula for rounded values. Find the doubling time of the bacteria. Round the doubling time to the nearest 0.01 hours.

OpenStudy (wolf1728):

For each hour, the population increases by 10,000 / 9,000 or 1.111111111111111 Ending Amount = Beginning Amount * (1+rate)^time Ending Amount = 9,000 * (1.11111111111111)^hours Ending Amount = 9,000 * (1.11111111111111)^1 = 10,000 Ending Amount = 9,000 * (1.11111111111111)^2 = 11,111 To find the doubling time, we must make the equation: 18,000 = 9,000 * (1.11111111111111)^time 2 = (1.11111111111111)^time taking logs of both sides log(2) = time * log (1.1111111111) time = log(2) / log (1.1111111111) time = 0.3010299957 / 0.0457574906 time = 6.5788134796 hours and to double check the calculations here's an online calculator: http://www.1728.org/expgrwth.htm

OpenStudy (anonymous):

thanks a lot!